OpenStudy (anonymous):

A quarterback claims that he can throw the football a horizontal distance of 196.6 m (215 yd). Furthermore, he claims that he can do this by launching the ball at the relatively low angle of 35° above the horizontal. To evaluate his claim, determine the speed with which this quarterback must throw the ball. Assume that the ball is launched and caught at the same vertical level and that air resistance can be ignored. For comparison, a baseball pitcher who can accurately throw a fastball at 45 m/s (100 mph) would be considered exceptional.

OpenStudy (anonymous):

This is an example of projectile motion. Use the formula Range=\[v^{2}\sin 2\theta/(g)\] where theta is angle of projection , v is velocity of projection, g is gravitational acceleration due to earth

OpenStudy (anonymous):

so this would look like V^2= sin(2*35)/9.8

OpenStudy (anonymous):

No 196.6 = v^2* sin(2*35)/9.8

OpenStudy (anonymous):

oooooh okay

OpenStudy (anonymous):

is this 196.6=(v^2/g)*[sin2*35/g]) or just V^2*sin92*35/g

OpenStudy (anonymous):

SANK YOUUUUU YOU SOOO KIND :)